(5 pts) 1. Select the pH of an aqueous solution in which the [OH-] = 5.00X10-?M. A) 0.301 B) 2.00 X 108 C) 7.70 D) 8.30
Answer #1

We know

pOH = - log [OH^-]

pOH = - log (5.00 * 10^-7) = - log 5.00 + 7 log 10

pOH = - 0.699 + 7 * 1 = 6.301

pH = 14.00 - pOH = 14.00 - 6.301 = 7.699

pH = 7.699 ~ 7.70

pH = 7.70